OpenStudy (anonymous):

tan^2x=(root2/2)sec x

OpenStudy (loser66):

tan^2 = sec^2 -1 so that it becomes sec ^2 -1(-sqrt(2))/2 sec =0 let t = sec then solve t^2 -(sqrt 2/2) t -1 =0